A flower pot is accidentally dropped from a building window 400m from the ground. Assuming it starts from rest, a How long will it take the pot to reach the ground?


1
Expert's answer
2022-03-10T18:01:15-0500

Explanations & Calculations


  • Initial velocity is given, the displacement it would make is given, and the acceleration is the gravitational acceleration.
  • So, with these apply "\small s =ut+1/2 at^2" for the pot's fall downwards.

"\qquad\qquad
\begin{aligned}
\small 400\,m&=\small 0t+\frac{1}{2}\times (+9.8\,ms^{-2})\times t^2\\
\small t^2&=\small \frac{400m\times2}{9.8\,ms^{-2}} \\
&=\small 81.6\,s 
\end{aligned}"